OpenStudy (anonymous):

what is dy/dx of y=tan squared x or y=tan^2 x

OpenStudy (anonymous):

using the chain rule let tanx = u, then y = u^2 du/dx = sec^2x and dy/du = 2u dy/dx = dy/du * du/dx = 2 tanx sec^2x

OpenStudy (anonymous):

simple (tanx)^2 derivate the outside....you will get the 2 as a coefficient and subtract 1 from the previous power..to get 2tanx....derivate the inside you will get sec^2x...multiply the two....to get dy/dx=2tanxsec^2x